Determine weather the following three points make a right triangle.
(2,-5), (-5,2), (-6,-3)

Answer by ewatrrr(24785)

 
Hi
(2, -5)
(-6,-3) m = -2/8 = -1/4
(-5, 2) m = -5/-1 = 5
Slopes of these segments indicate this is NOT a right triangle